PURE ROTATIONAL SPECTRUM OF THE TRANSITION METAL SILICIDE PtSi

\underline{DALE J. BRUGH} and RICHARD D. SUENRAM, National Institute of Standards and Technology, Optical Technology Division, Gaithersburg, MD 20899.

Laser vaporization of a Pt target rod in the throat of a supersonic expansion of Ar seeded with 1% SiH_{4} has produced the diatomic PtSi in sufficient abundance to allow the J=1-0 and J=2-1 pure rotational transitions in the v = 0 level of ^{194}Pt^{28}Si, ^{195}Pt^{28}Si, and ^{196}Pt^{28}Si to be recorded in the ^{1}Σ^{+} ground state. The measured line positions have allowed the rotational constant for each of these isotopic modifications to be determined as B(^{194}Pt^{28}Si)=4854.3637 MHz, B(^{195}Pt^{28}Si)=4851.2175 MHz, and B(^{196}Pt^{28}Si)=4848.1152 MHz. Work is currently under way to determine the dipole moment of PtSi.